OpenStudy (anonymous):

Find a Cartesian equation for the curve given by the polar equation: r+6cos(theta)=0

OpenStudy (caozeyuan):

\[r ^{2}+6r \cos \theta = 0\]

OpenStudy (caozeyuan):

\[x ^{2}+y ^{2}+6x=0\]

OpenStudy (anonymous):

it says the answer is (x+3)^2 + Y^2 = 9 can you take me through the process?

OpenStudy (caozeyuan):

After my previous work is done, this should be obvious

OpenStudy (caozeyuan):

DO you know a process called " Completing the Square " Your teacher should have taught you that

OpenStudy (caozeyuan):

plus 9 on BOTH sides. x^2+y^2+6x+9=9

OpenStudy (caozeyuan):

notice that x^2+6x+9 is just x+3)^2, so (x+3)^2+y^2=9 but your teacher should accept my answer

OpenStudy (caozeyuan):

@mlimsiaco12